# Gravity from entropy

### v1

@@ Abstract
In this Letter, a continuum modified theory of gravity based on a statistical mechanics action is considered. This theory treats the metric at each point of spacetime as a density matrix. Specifically, this work proposes an action based on the quantum entropy between the metric and the metric induced by the matter fields. A crucial aspect of the presented theory is the adoption of a topological (Dirac-Kälher like [22, 23]) description of bosonic matter fields. Note that the extension of Dirac-Kälher and staggered fermions formalism to bosonic particles is gaining increasing interest in lattice gauge theory [24, 25]. and network theory as well [26]. These bosonic matter fields are described as the direct sum of a zero-form and a one-form defined on the Riemannian manifold describing spacetime. Moreover the metric induced by the topological matter fields is expressed in terms of the Hodge-Dirac operator [27]. This work greatly expands on previous results [26] obtained in the discrete setting by the same author. On one side, here a continuum and fully Lorentz invariant theory is proposed. This progress is based on the development of the suitable mathematical framework to define the Lorentz invariant entropy and cross-entropy between the metric of the spacetime and the metric induced by the matter fields. On the other side, here the relation of this statistical mechanics/information theory action with the Einstein-Hilbert action [28, 29] is established defining a clear connection to gravity. Indeed by considering a continuum limit, in this Letter the intrinsic difficulty related to the definition of the curvature of networks, simplicial and cell complexes is avoided.
@@
In order to derive gravity from our entropic action we need to consider the topological bosonic matter field. The topological bosonic matter field is a type of Dirac-Khäler [22, 23] boson given by the direct sum of a zero-form and a one-form. Topological bosonic fields are receiving increasing attention in discrete theories developed in network [26] and lattice gauge theories [24, 25]. Taking into consideration topological bosonic fields will allow us to introduce in the metric induced by the matter fields, terms depending on the mass of the bosonic field. Thus, in this way, we address the first limitation of the warm-up scenario that we have presented above.
@@
Indicating here with ${d}$ the differential operator and with ${\delta}$ the codifferential operator, we define the Dirac operator $D$ as
@@
This action defines a modified theory of gravity. For $\alpha\ll 1$ and $\beta\ll 1$ this action reduces up to a constant, to the Einstein-Hilbert action coupled with the scalar topological field. Indeed in this limit $\bm{\mathcal{L}}$ reduces to
@@
where $\Lambda^{\prime}$ is the rescaled cosmological constant given by $\Lambda^{\prime}=(\Lambda+(d+1))/4\beta.$ Thus $\Lambda^{\prime}>0$ also if $\Lambda=0$. In the interesting limit in which $\omega_{\alpha}=0$ and the bosonic matter field is only scalar, Eq.(39) reduces (up to a constant) to the widely studied Einstein-Hilbert action [28, 29] coupled with the scalar field [33],
@@
In conclusion this work proposes a modified theory of gravity emerging from statistical mechanics and information theory action. The approach is based on the interpretation of the metric as a density matrix albeit its trace is not unitary. The statistical mechanics action is given by the quantum relative entropy between the metric and the metric induced by the matter fields. Here a Lorentzian theory consistent with this fundamental physical interpretation of gravity if formulated. In order to do this we have built the necessary mathematical background to define the entropy and the cross-entropy of rank 2 tensors in a Lorentzian well defined way. A modified gravity is emerging from this framework when the matter fields are described by topological Dirac-Kälher bosons formed by the direct sum between a zero and a one form and the induced metric also depends on the curvature of spacetime. The modified Einstein equations reduce to the Einstein equations in the regime of low coupling.

### v2

@@ Abstract
In this Letter, a continuum modified theory of gravity based on a statistical mechanics action is considered. This theory treats the metric at each point of spacetime as a density matrix. Specifically, this work proposes an action based on the quantum entropy between the metric and the metric induced by the matter fields. A crucial aspect of the presented theory is the adoption of a topological (Dirac-Kähler like [22, 23]) description of bosonic matter fields. Note that the extension of Dirac-Kähler and staggered fermions formalism to bosonic particles is gaining increasing interest in lattice gauge theory [24, 25]. and network theory as well [26]. These bosonic matter fields are described as the direct sum of a zero-form and a one-form defined on the Riemannian manifold describing spacetime. Moreover the metric induced by the topological matter fields is expressed in terms of the Hodge-Dirac operator [27]. This work greatly expands on previous results [26] obtained in the discrete setting by the same author. On one side, here a continuum and fully Lorentz invariant theory is proposed. This progress is based on the development of the suitable mathematical framework to define the Lorentz invariant entropy and cross-entropy between the metric of the spacetime and the metric induced by the matter fields. On the other side, here the relation of this statistical mechanics/information theory action with the Einstein-Hilbert action [28, 29] is established defining a clear connection to gravity. Indeed by considering a continuum limit, in this Letter the intrinsic difficulty related to the definition of the curvature of networks, simplicial and cell complexes is avoided.
@@
In order to derive gravity from our entropic action we need to consider the topological bosonic matter field. The topological bosonic matter field is a type of Dirac-Kähler [22, 23] boson given by the direct sum of a zero-form and a one-form. Topological bosonic fields are receiving increasing attention in discrete theories developed in network [26] and lattice gauge theories [24, 25]. Taking into consideration topological bosonic fields will allow us to introduce in the metric induced by the matter fields, terms depending on the mass of the bosonic field. Thus, in this way, we address the first limitation of the warm-up scenario that we have presented above.
@@
Indicating here with ${d}\phi$ the differential operator applied to $\phi$ and with ${\delta}\omega$ the codifferential operator applied to $\omega$, we define the Dirac operator $D$ as
@@
This action defines a modified theory of gravity. For $\alpha\ll 1$ and $\beta\ll 1$ this action reduces to the Einstein-Hilbert action coupled with the scalar topological field. Indeed in this limit $\bm{\mathcal{L}}$ reduces to
@@
where $\Lambda^{\prime}$ is the rescaled cosmological constant given by $\Lambda^{\prime}=\Lambda/(4\beta).$ In the interesting limit in which $\omega_{\alpha}=0$ and the bosonic matter field is only scalar, Eq.(39) reduces to the widely studied Einstein-Hilbert action [28, 29] coupled with the scalar field [33],
@@
In conclusion this work proposes a modified theory of gravity emerging from statistical mechanics and information theory action. The approach is based on the interpretation of the metric as a density matrix albeit its trace is not unitary. The statistical mechanics action is given by the quantum relative entropy between the metric and the metric induced by the matter fields. Here a Lorentzian theory consistent with this fundamental physical interpretation of gravity if formulated. In order to do this we have built the necessary mathematical background to define the entropy and the cross-entropy of rank 2 tensors in a Lorentzian well defined way. A modified gravity is emerging from this framework when the matter fields are described by topological Dirac-Kähler bosons formed by the direct sum between a zero and a one form and the induced metric also depends on the curvature of spacetime. The modified Einstein equations reduce to the Einstein equations in the regime of low coupling.